# Breather interactions in the integrable discrete Manakov system and trigonometric Yang-Baxter maps

### v1

@@ 3 Transmission coefficients of soliton and breather solutions
Taking the limit as $n\to-\infty$ in the above equation, and considering that ${\mathbf{M}}_{n}^{(\text{up})}(z)\to{\mathbf{I}}_{2}$ and ${\mathbf{M}}_{n}^{(\text{up})}(z)\to{\mathbf{0}}_{2}$ one obtains the first of (3.1), which can then be extended into $D^{-}$ by analytical continuation. A similar approach works for ${\mathbf{c}}(z)$. The coefficients ${\mathbf{a}}(z),\,\bar{{\mathbf{a}}}(z)$ can be found via similar limits of the “left” eigenfunctions, or directly through the symmetries in (2.11a).
@@
Note that the above expressions correspond to the second case in Proposition 1, i.e., each of the discrete eigenvalue in the octet is a double zero of the determinant of one of the coefficients, all of which have rank 0 (i.e., they vanish) when evaluated at the eigenvalues in the pertinent region of analyticity.
@@
where ${\mathbf{u}}_{1}=\boldsymbol{\gamma}_{1}^{*}/\|\boldsymbol{\gamma}_{1}\|$ and ${\mathbf{v}}_{1}=\boldsymbol{\delta}_{1}^{*}/\|\boldsymbol{\delta}_{1}\|$ are normalized polarization vectors. Note that when ${\mathbf{v}}_{1}=(1,0)^{T}$, $\bar{{\mathbf{c}}}(z)$ in (3.3a) and ${\mathbf{a}}(z)$ in (3.3c) reduce to diagonal matrices while ${\mathbf{c}}(z)$ in (3.3b) and $\bar{{\mathbf{a}}}(z)$ in (3.3d) remain nontrivial, in agreement with the result obtained in [4] for the fundamental soliton case. The expressions of the transmission coefficients for fundamental and composite breathers had not been determined before.
@@ 4.1 Generic multi-soliton solution
[…]
@@ 4.2 Fundamental soliton and breather interactions
from which it can be seen that $e^{-\zeta_{2}}\rightarrow 0$ as $\tau\rightarrow-\infty$. Thus, in this limit, all terms in the linear system that contain a norming constant labeled “2” can be neglected, in light of (4.8) and (4.9). The solution of the resulting asymptotic system in this case is straightforward, and it is given in detail in Appendix B. The result is that as $\tau\rightarrow-\infty$ with $\zeta_{1}$ fixed, the 2-fundamental breather solution approaches a 1-fundamental breather identical to (2.29a). Similarly, in the reference frame of fundamental breather “2”, one can express
@@
From this, $e^{-\zeta_{1}}\rightarrow 0$ as $\tau\rightarrow+\infty$, in which case all terms in the linear system that contain a norming constant labeled “1” can be neglected. In this limit, one finds that the 2-fundamental breather solution again approaches a form identical to (2.29a), but with parameters labeled “2”. A conclusion can then be drawn that the prescribed polarization vectors $\boldsymbol{\gamma}_{j}$ and $\boldsymbol{\delta}_{j}$ can be attributed to long time limits in the following way:
@@ 4.3 Explicit polarization maps
Then, the above can be substituted into (4.12a) and (4.12c) so that all “$+$” polarization vectors are expressed in terms of “--” ones. After simplification, we have:
@@
Note that from the form given in (4.19), we see that the shifts are symmetric with respect to interchanging $1\leftrightarrow 2$ and +$\leftrightarrow$-+ ↔ -. On the other hand, (4.24) are no longer symmetric upon interchanging $1\leftrightarrow 2$, consistently with what one observes in the Manakov formulas (1.4).
@@
As long as $\mu$ and $\kappa$ are both nonzero, both components of $\boldsymbol{\delta}_{1}^{+}$ from the above equation are nonzero, so soliton 1 becomes itself a fundamental breather upon interaction with the fundamental breather. This phenomenon can be observed in Fig. 2.
@@ 4.4 Interactions involving composite breathers
Let us assume soliton 1 is either a fundamental soliton or a fundamental breather (i.e., with a rank-1 norming constant), while soliton 2 is a composite breather, i.e., we take:
@@
where $\boldsymbol{\delta}_{2}=(1,0)^{T}$, $\tilde{\boldsymbol{\delta}}_{2}=(0,1)^{T}$, and $\boldsymbol{\gamma}_{2}$ and $\boldsymbol{\varepsilon}_{2}$ are linearly independent vectors. From Eqs. (4.12) we have:
@@
and since the for the composite breather:
@@
showing that the rank-1 soliton (be it a fundamental soliton or a fundamental breather) is essentially unaffected by the interaction with the composite breather, as the interaction only results in a shift in the overall phase and in the soliton center. For the composite breather, assuming again that like in the ccSPE $\boldsymbol{\varepsilon}_{2}$ transforms like $\boldsymbol{\gamma}_{2}$, and $\tilde{\boldsymbol{\delta}}_{2}$ transforms like $\boldsymbol{\delta}_{2}$ in (4.12), with the appropriate transmission coefficients, we find:
@@
Figure 7: The same interaction as in Fig. 6, with the predicted asymptotic states subtracted in each direction. In particular; in (a) the fundamental soliton with polarization vectors $\boldsymbol{\gamma}_{1}^{-},\boldsymbol{\delta}_{1}^{-}$ is subtracted, in (b) the fundamental soliton with polarization vectors $\boldsymbol{\gamma}_{1}^{+},\boldsymbol{\delta}_{1}^{+}$ is subtracted, in (c) the composite breather with polarization vectors $\boldsymbol{\gamma}_{2}^{-},\boldsymbol{\varepsilon}_{2}^{-},\boldsymbol{\delta}_{2}^{-},\tilde{\boldsymbol{\delta}}_{2}^{-}$ is subtracted, in (d) the composite breather with polarization vectors $\boldsymbol{\gamma}_{2}^{+},\boldsymbol{\varepsilon}_{2}^{+},\boldsymbol{\delta}_{2}^{+},\tilde{\boldsymbol{\delta}}_{2}^{+}$ is subtracted.
@@ 5 Refactorization problem and Yang-Baxter maps
In this section, we reinterpret the two-body interactions in the discrete Manakov model in terms of Yang-Baxter maps by recasting the condition on the scattering coefficients that leads to the maps for the polarization vectors before and after the interaction as a refactorization problem.
@@ 5.1 Generalities
We recall some general facts and notions from the theory of refactorization and Yang-Baxter maps, following e.g. [32, 24]. Let $A({\mathbf{x}},\lambda)$ be a given matrix-valued function that depends on a point ${\mathbf{x}}\in X$, $X$ being some set (typically $\mathbb{C}\mathbb{P}^{n}$ or a Grassmannian in the context of soliton interactions), a (spectral) parameter $\lambda\in\mathbb{C}$, and possibly other (model-dependent) parameters. Consider the refactorization problem
@@ 5.2 Application to interactions in the integrable discrete Manakov model
The point is that the two-body interactions in the discrete Manakov model can be cast into the (reduced) refactorization problem (5.8) where the role of the simple elements is played by the scattering coefficients in (3.2) and (3.3). Given that the for composite breathers the transmission coefficients (3.2) are proportional to the identity, the refactorization is trivial and no interesting Yang-Baxter map arises. Thus we focus on the fundamental breather coefficients (the fundamental soliton being a special case) and restrict our attention to (3.3b) since the structure of (3.3a) is similar. Specifically, consider the coefficients ${\mathbf{c}}_{j}(z,\boldsymbol{\gamma}_{j})$ in (4.14) and the refactorization
@@
Third, the projectors here are rank one Hermitian projectors in $\mathbb{C}^{2}$ and the correspondence with (5.9) is $\alpha_{j}\to\lambda_{j}$, $\Pi_{j}\to\Pi_{j}^{-}=\frac{\boldsymbol{\gamma}^{-}_{j}\boldsymbol{\gamma}_{j}^{-\dagger}}{\boldsymbol{\gamma}_{j}^{-\dagger}\boldsymbol{\gamma}^{-}_{j}}$, $P_{j}\to\Pi_{j}^{+}=\frac{\boldsymbol{\gamma}^{+}_{j}\boldsymbol{\gamma}_{j}^{+\dagger}}{\boldsymbol{\gamma}_{j}^{+\dagger}\boldsymbol{\gamma}_{j}}$, $j=1,2$. Therefore, we have a parametric Yang-Baxter map on rank one Hermitian projectors of trigonometric type, as seen from the reparametrization (5.11).
@@
$=\left(r^{2}+(s_{1}-s_{2})^{2}+4s_{1}s_{2}\frac{|\gamma_{1}^{-\dagger}\gamma_{2}^{-}|^{2}}{||\gamma_{1}^{-}||^{2}||\gamma_{2}^{-}||^{2}}\right)\Pi^{-}_{j}$
@@
$||\phi\boldsymbol{\gamma}_{j}^{-}||^{2}=\boldsymbol{\gamma}_{j}^{-\dagger}\,\phi^{\dagger}\phi\,\boldsymbol{\gamma}^{-}_{j}=\Delta^{2}\,||\boldsymbol{\gamma}_{j}^{-}||^{2}\,.$ (5.17)
@@
$\frac{\boldsymbol{\gamma}_{j}^{+}\boldsymbol{\gamma}_{j}^{+\dagger}}{||\boldsymbol{\gamma}_{j}^{+}||^{2}}=\Pi_{j}^{+}=\phi\Pi_{j}^{-}\phi^{-1}=\frac{1}{\Delta^{2}}\phi\Pi_{j}^{-}\phi^{\dagger}=\frac{\phi\boldsymbol{\gamma}_{j}^{-}(\phi\boldsymbol{\gamma}_{j}^{-})^{\dagger}}{||\phi\boldsymbol{\gamma}_{j}^{-}||^{2}}\,.$ (5.18)
@@
which we can interpret as a map between elements $[\boldsymbol{\gamma}_{j}^{\pm}]$ in $\mathbb{C}\mathbb{P}^{1}$ given in terms of representatives $\boldsymbol{\gamma}_{j}^{\pm}$ in $\mathbb{C}^{2}$, if one is not interested in normalizations. Of course, physically it is important to determine the normalizations precisely, and the corresponding maps are then then ones given in (4.24a)-(4.24b).
@@ 5.3 Common structures and differences across three integrable models
[…]
@@
[…]
@@ 6 Concluding remarks
In this work, we presented a complete characterization of the various soliton and breather interactions in the integrable discrete Manakov model. This was done using a generalization of the Manakov method [which essentially expresses the change in the polarization of the soliton/breather after the interaction via the transmission coefficients associated to the interacting soliton/breather] to include fundamental and composite breathers. As is well-known, interactions in multicomponent integrable PDEs (and, as shown here, in their integrable discretizations) are intimately related to set-theoretical solutions of the Yang-Baxter equation with spectral parameters, or parametric Yang-Baxter maps. In particular, the explicit formulas for the transmission coefficients that characterize fundamental solitons and fundamental breathers allowed us to interpret the interactions in terms of a refactorization problem, and derive the associated Yang-Baxter maps describing the effect of interactions on the soliton polarizations. As an essential feature of the integrable discrete Manakov model, a novel Yang–Baxter map of trigonometric type was derived.
A natural follow-up question concerns the type of set-theoretical solutions of the reflection equation, or reflection maps [14, 12], that could be constructed from the trigonometric Yang-Baxter map found here. This is particularly relevant as a fruitful method to construct such reflection maps is to pose the integrable model on the half-line (or positive integers) and study the corresponding integrable boundary conditions. Since the Ablowitz-Ladik with such boundary conditions has been thoroughly studied [9, 10, 11], we expect that an extension to the discrete Manakov model would yield reflection maps of trigonometric type. In turn, this is of relevance to the study of the set-theoretical Yang-Baxter and reflection equations in the context of (skew)-braces and other related algebraic structures. Recently, the importance of developing parametric versions of these structures has emerged in [15, 16]. This is left for future work.

### v2

@@ 3 Transmission coefficients of soliton and breather solutions
Taking the limit as $n\to$-$\infty$ in the above equation, and considering that ${\mathbf{M}}_{n}^{(\text{up})}(z)\to{\mathbf{I}}_{2}$ and $\bar{{\mathbf{M}}}_{n}^{(\text{up})}(z)\to{\mathbf{0}}_{2}$ one obtains the first of (3.1), which can then be extended into $D^{-}$ by analytic continuation. A similar approach works for ${\mathbf{c}}(z)$. The coefficients ${\mathbf{a}}(z),\,\bar{{\mathbf{a}}}(z)$ can be found via similar limits of the “left” eigenfunctions, or directly through the symmetries in (2.11a).
@@
Note that the above expressions correspond to the second case in Proposition 1, i.e., each of the discrete eigenvalues in the octet is a double zero of the determinant of one of the coefficients, all of which have rank 0 (i.e., they vanish) when evaluated at the eigenvalues in the pertinent region of analyticity.
@@
where ${\mathbf{u}}_{1}=\boldsymbol{\gamma}_{1}^{*}/\|\boldsymbol{\gamma}_{1}\|$ and ${\mathbf{v}}_{1}=\boldsymbol{\delta}_{1}^{*}/\|\boldsymbol{\delta}_{1}\|$ are normalized polarization vectors. Note that when ${\mathbf{v}}_{1}=(1,0)^{T}$, $\bar{{\mathbf{c}}}(z)$ in (3.3a) and ${\mathbf{a}}(z)$ in (3.3c) reduce to diagonal matrices while ${\mathbf{c}}(z)$ in (3.3b) and $\bar{{\mathbf{a}}}(z)$ in (3.3d) remain nontrivial, in agreement with the result obtained in [4] for the fundamental soliton case. The expressions of the transmission coefficients for fundamental and composite breathers were not determined before.
@@ 4.1 Generic multi-soliton solution
$\boldsymbol{\chi}\equiv(\boldsymbol{\chi}_{1},\boldsymbol{\chi}_{2},\dots,\boldsymbol{\chi}_{2J})=\left(\bar{{\mathbf{N}}}^{\prime\,\textrm{(up)}}_{n}(\bar{z}_{1}),\dots,\bar{{\mathbf{N}}}^{\prime\,\textrm{(up)}}_{n}(\bar{z}_{J}),\bar{{\mathbf{N}}}^{\prime\,\textrm{(up)}}_{n}(i/z_{1}),\dots,\bar{{\mathbf{N}}}^{\prime\,\textrm{(up)}}_{n}(i/z_{J})\right).$ (4.1b)
@@ 4.2 Fundamental soliton and breather interactions
from which it can be seen that $e^{-\zeta_{2}}\rightarrow 0$ as $\tau\rightarrow$-$\infty$. Thus, in this limit, all terms in the linear system that contain a norming constant labeled “2” can be neglected, in light of (4.8) and (4.9). The solution of the resulting asymptotic system in this case is straightforward, and it is given in detail in Appendix B. The result is that as $\tau\rightarrow$-$\infty$ with $\zeta_{1}$ fixed, the 2-fundamental breather solution approaches a 1-fundamental breather identical to (2.29). Similarly, in the reference frame of fundamental breather “2”, one can express
@@
From this, $e^{-\zeta_{1}}\rightarrow 0$ as $\tau\rightarrow+\infty$, in which case all terms in the linear system that contain a norming constant labeled “1” can be neglected. In this limit, one finds that the 2-fundamental breather solution again approaches a form identical to (2.29), but with parameters labeled “2”. A conclusion can then be drawn that the prescribed polarization vectors $\boldsymbol{\gamma}_{j}$ and $\boldsymbol{\delta}_{j}$ can be attributed to long time limits in the following way:
@@ 4.3 Explicit polarization maps
Then, the above can be substituted into (4.12a) and (4.12c) so that all “++” polarization vectors are expressed in terms of “-” ones. After simplification, we have:
@@
Note that from the form given in (4.19), we see that the shifts are symmetric with respect to interchanging $1\leftrightarrow 2$ and $\leftrightarrow-$. On the other hand, (4.24) are no longer symmetric upon interchanging $1\leftrightarrow 2$, consistent with what one observes in the Manakov formulas (1.4).
@@
If $\mu$ and $\kappa$ are both nonzero, both components of ${\mathbf{v}}_{1}^{+}$ from the above equation are generically nonzero, so soliton 1 becomes itself a fundamental breather upon interaction with the fundamental breather. This phenomenon can be observed in Fig. 2.
@@ 4.4 Interactions involving composite breathers
Let us assume soliton 1 is either a fundamental soliton or a fundamental breather (i.e., with a rank-1 norming constant), while soliton 2 is a composite breather, i.e., we take
@@
where $\boldsymbol{\delta}_{2}=(1,0)^{T}$, $\tilde{\boldsymbol{\delta}}_{2}=(0,1)^{T}$, and $\boldsymbol{\gamma}_{2}$ and $\boldsymbol{\varepsilon}_{2}$ are linearly independent vectors. From Eqs. (4.12) we have
@@
and since for the composite breather
@@
showing that the rank-1 soliton (be it a fundamental soliton or a fundamental breather) is essentially unaffected by the interaction with the composite breather, as the interaction only results in a shift in the overall phase and in the soliton center. For the composite breather, assuming again that like in the ccSPE $\boldsymbol{\varepsilon}_{2}$ transforms like $\boldsymbol{\gamma}_{2}$, and $\tilde{\boldsymbol{\delta}}_{2}$ transforms like $\boldsymbol{\delta}_{2}$ in (4.12), with the appropriate transmission coefficients, we find
@@
Figure 7: The same interaction as in Fig. 6, with the predicted asymptotic states subtracted in each direction. In particular, in (a) the fundamental soliton with polarization vectors $\boldsymbol{\gamma}_{1}^{-},\boldsymbol{\delta}_{1}^{-}$ is subtracted, in (b) the fundamental soliton with polarization vectors $\boldsymbol{\gamma}_{1}^{+},\boldsymbol{\delta}_{1}^{+}$ is subtracted, in (c) the composite breather with polarization vectors $\boldsymbol{\gamma}_{2}^{-},\boldsymbol{\varepsilon}_{2}^{-},\boldsymbol{\delta}_{2}^{-},\tilde{\boldsymbol{\delta}}_{2}^{-}$ is subtracted, in (d) the composite breather with polarization vectors $\boldsymbol{\gamma}_{2}^{+},\boldsymbol{\varepsilon}_{2}^{+},\boldsymbol{\delta}_{2}^{+},\tilde{\boldsymbol{\delta}}_{2}^{+}$ is subtracted.
@@ 5 Refactorization problem and Yang-Baxter maps
In an interaction of $J$ fundamental breathers with velocities $v_{1}<v_{2}<\cdots<v_{J}$, the net change in the polarization of each soliton from $\tau\rightarrow$-$\infty$ to $\tau\rightarrow+\infty$ is independent of the order in which the intermediate interactions occur. As was established in [7] in the special case of fundamental solitons, this result is tied to the fact that the polarization map (4.24) associated with each two-body interaction is a Yang-Baxter map. In this section, we reinterpret the two-body interactions in the discrete Manakov model in terms of Yang-Baxter maps by recasting the condition on the scattering coefficients that leads to the maps for the polarization vectors before and after the interaction as a refactorization problem.
@@ 5.1 Generalities
We recall some general facts and notions from the theory of refactorization and Yang-Baxter maps, following e.g. [32, 24]. Let $A({\mathbf{x}},\lambda)$ be a given matrix-valued function that depends on a point ${\mathbf{x}}\in X$, with $X$ being some set (typically $\mathbb{C}\mathbb{P}^{N}$ or a Grassmannian in the context of soliton interactions), a (spectral) parameter $\lambda\in\mathbb{C}$, and possibly other (model-dependent) parameters. Consider the refactorization problem
@@ 5.2 Application to interactions in the integrable discrete Manakov model
The point is that the two-body interactions in the discrete Manakov model can be cast into the (reduced) refactorization problem (5.8) where the role of the simple elements is played by the scattering coefficients in (3.2) and (3.3). Given that for composite breathers the transmission coefficients (3.2) are proportional to the identity, the refactorization is trivial and no interesting Yang-Baxter map arises. Thus we focus on the fundamental breather coefficients (the fundamental soliton being a special case) and restrict our attention to (3.3b) since the structure of (3.3a) is similar. Specifically, consider the coefficients ${\mathbf{c}}_{j}(z,\boldsymbol{\gamma}_{j})$ in (4.14) and the refactorization
@@
Third, the projectors here are rank one Hermitian projectors in $\mathbb{C}^{2}$ and the correspondence with (5.9) is $\alpha_{j}\to\lambda_{j}$, $\Pi_{j}\to\Pi_{j}^{-}=\frac{\boldsymbol{\gamma}^{-}_{j}\boldsymbol{\gamma}_{j}^{-\dagger}}{\boldsymbol{\gamma}_{j}^{-\dagger}\boldsymbol{\gamma}^{-}_{j}}$, $P_{j}\to\Pi_{j}^{+}=\frac{\boldsymbol{\gamma}^{+}_{j}\boldsymbol{\gamma}_{j}^{+\dagger}}{\boldsymbol{\gamma}_{j}^{+\dagger}\boldsymbol{\gamma}_{j}}$, $j=1,2$. Therefore, we have a parametric Yang-Baxter map on rank one Hermitian projectors of trigonometric type, as seen from the reparametrization (5.11). Strictly from the point of view of the YB maps, there is no essential new effect on the polarization shifts from this trigonometric map as compared for instance to the rational. Indeed, the last reparametrization (5.12) shows that in terms of $\lambda$, the YB map takes the same form as in the rational case. However, physically speaking, one has to remember that the trigonometric nature of the map is tied into the similar trigonometric nature of the Lax matrices in (2.1). In turn, this dictates how the physical properties of the solitons such as their velocity and amplitudes are related to the spectral parameters (see e.g. (2.27a)-(2.27c)) and therefore how they influence the polarization shifts. Given the same velocity say of a soliton in the rational case and the trigonometric case, the corresponding effect on the polarization shifts when interacting with another soliton will be different in the two cases.
@@
[…]
@@
[…]
@@
[…]
@@
which we can interpret as a map between elements $\boldsymbol{\gamma}_{j}^{\pm}]$ in $\mathbb{C}\mathbb{P}^{1}$ given in terms of representatives $\boldsymbol{\gamma}_{j}^{\pm}$ in $\mathbb{C}^{2}$, if one is not interested in normalizations. Of course, physically it is important to determine the normalizations precisely, and the corresponding maps are then the ones given in (4.24a)-(4.24b).
@@ 5.3 Common structures and differences across three integrable models
$\beta_{j}(k)=\frac{k_{j}^{*}-k_{j}}{k-k_{j}^{*}}~~\text{(vector NLS)}\,$, (5.21)
$\beta_{j}(k)=\frac{k^{2}}{k_{j}^{2}}\frac{k_{j}^{*2}-k_{j}^{2}}{k^{2}-k_{j}^{*2}}~~\text{(ccSPE)}\,$, (5.22)
$\beta_{j}(k)=\frac{(\bar{k}_{j}^{2}-k_{j}^{2})(\bar{k}_{j}^{2}-k_{j}^{-2})}{(k^{2}-\bar{k}_{j}^{2})(k^{-2}+\bar{k}_{j}^{2})}~~\text{(discrete Manakov)}\,.$ (5.23)
@@
$\beta_{j}(k)=\frac{\lambda-\lambda_{j}}{\lambda-\lambda_{j}^{*}}-1\,,~~k\to\lambda~~\text{(vector NLS)}\,$, (5.25)
$\beta_{j}(k)=\frac{\frac{1}{k^{2}}-\frac{1}{k_{j}^{2}}}{\frac{1}{k^{2}}-\frac{1}{k_{j}^{*2}}}-1=\frac{\lambda-\lambda_{j}}{\lambda-\lambda_{j}^{*}}-1\,,~~1/k^{2}\to\lambda~~\text{(ccSPE)}\,.$ (5.26)
@@ 6 Concluding remarks
In this work, we presented a complete characterization of the various soliton and breather interactions in the integrable discrete Manakov model. This was done using a generalization of the Manakov method (which essentially expresses the change in the polarization of the soliton/breather after the interaction via the transmission coefficients associated to the interacting soliton/breather) to include fundamental and composite breathers. As is well-known, interactions in multicomponent integrable PDEs (and, as shown here, in their integrable discretizations) are intimately related to set-theoretical solutions of the Yang-Baxter equation with spectral parameters, or parametric Yang-Baxter maps. In particular, the explicit formulas for the transmission coefficients that characterize fundamental solitons and fundamental breathers allowed us to interpret the interactions in terms of a refactorization problem, and derive the associated Yang-Baxter maps describing the effect of interactions on the soliton polarizations. As an essential feature of the integrable discrete Manakov model, a novel Yang–Baxter map of trigonometric type was derived. In that respect, we expect these maps to possess nice Poisson and symplectic properties arising from the underlying Hamiltonian structure of the discrete Manakov model. The latter would be the natural vector generalization of the one presented e.g. in [11] which relied on the trigonometric classical $r$-matrix in the sense of the Belavin-Drinfeld classification [8]. Such nice properties were studied in detail in [25, 26] in relation to the rational case. The investigation of these properties in the trigonometric case is beyond the scope of this paper, but we expect it will be the subject of future work.
Another natural follow-up question concerns the type of set-theoretical solutions of the reflection equation, or reflection maps [14, 12], that could be constructed from the trigonometric Yang-Baxter map found here. This is particularly relevant as a fruitful method to construct such reflection maps is to pose the integrable model on the half-line (or positive integers) and study the corresponding integrable boundary conditions. Since the Ablowitz-Ladik with such boundary conditions has been thoroughly studied [9, 10, 11], we expect that an extension to the discrete Manakov model would yield reflection maps of trigonometric type. In turn, this is of relevance to the study of the set-theoretical Yang-Baxter and reflection equations in the context of (skew)-braces and other related algebraic structures. Recently, the importance of developing parametric versions of these structures has emerged in [15, 16]. This is left for future work.
Finally, we remark that the 2-component vector system (1.1) has an integrable $N$-component generalization (see for example [31]), and a study similar to that of the present paper could be applied to this generic case. The norming constants would be $N\times N$ matrices, so one could consider the interactions of solitons associated with $N$ possible ranks. In general, this would produce YB maps relating to Grassmannians and the variety of possible interactions between solitons corresponding to different ranks of projectors will increase substantially.